OpenStudy (anonymous):

What is the square root of -23?

OpenStudy (anonymous):

that will be an imaginary number because it's not possible to have negative square roots.

OpenStudy (mathstudent55):

Have you learned imaginary numbers?

OpenStudy (anonymous):

No I haven't. I have to find the principal square root & the negative square root. Also I must determine if it is rational and/or a perfect square.

OpenStudy (mathstudent55):

If so, then \( \sqrt{-23} = \sqrt{-1 \times 23} = \sqrt{-1}\sqrt{23} = i\sqrt{23} \)

OpenStudy (mathstudent55):

If you haven't learned imaginary numbers, then the answer is that -23 has no square root.

OpenStudy (anonymous):

So, it is undefined?

OpenStudy (mathstudent55):

Yes.

OpenStudy (anonymous):

What about for -36? @mathstudent55